OpenStudy (anonymous):

2 4 % × 70 = _____ Select one of the options below as your answer: A. .168 B. .175 C. .28 D. .35

OpenStudy (anonymous):

2/4=0.5 0.5/100=0.005 0.005*70=0.35

OpenStudy (anonymous):

oh, thats a fraction

OpenStudy (anonymous):

yes you are correct then